Please help me solve this question: Write a point-slope equation for the line that passes through the point (11, 16) and is parallel to the line given by y = 9x + 41. Do not use parenthesis on the y side.
Answer by dragonwalker(73)

If the line is running parallel then we know that the values of x and y will be in relation to each other, but the integer will be different.
So we need to find how much this difference is.
We can do this by substituting y and x into the known equation and then we can solve the variable.
Let us call the variable z
So y = 9x + z
16 = 9(11) + z
16 = 99 + z
16 - 99 = z
so z = -83
Therefore the formula for the line will be:
y = 9x - 83
to check substitute value of x into the formula (11 )and y should be equal to 16:
y = 9(11) -83
y = 99 - 83
y = 16!!!